OpenStudy (anonymous):

The atomic mass of an element whose atoms consist of eight protons, nine neutrons, and eight electrons is ?

thomaster (thomaster):

number of protons = atomic number protons + neutrons = mass number

OpenStudy (frostbite):

\[m(element)=8*m(proton)+9*m(neutron)+8*m(electron)\] Assume just to make it more easy that: m(proton)=m(neutron)=1 u* m(electron)= 1/1000 u 1 u=*1,66053886 · 10^-27 kg

OpenStudy (frostbite):

It is a isotope remember that.

OpenStudy (frostbite):

beside the 15.9994 is the average mass of the natural occurring isotopes multiplied with their relative percentage of appearance :)

thomaster (thomaster):

them it's 16.99913170 amu :P

thomaster (thomaster):

anyway it would be something close to 17 as 8+9 = 17 :P